OpenStudy (anonymous):

Two positive integers have a sum of 10 and a product of 21. What are the integers?

OpenStudy (id21):

you need to solve a system x+y=10 and x*y=21

OpenStudy (anonymous):

How do you solve the system?

OpenStudy (id21):

\[x+y=10; x=10-y,\] you must put it in other equation\[x \times y = 21; (10-y)y=21; y^2-10y+21=0\] So \[y_{1,2}=7; 3\] and use it in x=10-y, so x=3 when y=7 and x-7 when y=3